OpenStudy (anonymous):

HELP!!! someone? Explain this to me please! Suppose a house that costs $270,000 appreciates by 5% each year. In about how many years will the house be worth $350,000? Use the equation 350 = (270)(1.05)x and round the value of x to the nearest year. 5 4 6 7

zepdrix (zepdrix):

So we have an exponential equation:\[\Large\rm A=P(b)^x\]Where P is the principle. (Starting amount). We assume the house will be worth 350 (thousand) at some point, so we let our A be that amount. b represents the growth rate.\[\Large\rm 350=270(1.05)^x\]To solve for x we'll need to use logarithms.

zepdrix (zepdrix):

Let's start by dividing each side by 270,\[\Large\rm \frac{350}{270}=1.05^x\]

OpenStudy (anonymous):

So itll be 1.29= 1.05^x

zepdrix (zepdrix):

1.30 is probably a better way to round that.

OpenStudy (anonymous):

True

zepdrix (zepdrix):

Next, we'll take the log of each side (log base 10 is fine),\[\Large\rm \log(1.30)=\log(1.05^x)\]

OpenStudy (anonymous):

0.1139=log(1.05^x)

zepdrix (zepdrix):

The log is helpful for getting variables out of the exponent position.

OpenStudy (anonymous):

Hmm but my calculator doesnt register x as a exponent

zepdrix (zepdrix):

Here is the log rule that will help us:\[\Large\rm \log(a^\color{orangered}{x})=\color{orangered}{x} \log(a)\]

zepdrix (zepdrix):

We're not up to the calculator step yet ;)

OpenStudy (anonymous):

Oh okay! so it would look like 0.1139=xlog(1.05)

zepdrix (zepdrix):

Mmm good good good. Just a few more steps to finish it up from there.

OpenStudy (anonymous):

Okays :D

zepdrix (zepdrix):

It seems like you prefer to carry the decimals around, rather than the logs. So from here, let's get our decimal approximation for log(1.05)

OpenStudy (anonymous):

its 0.0211

zepdrix (zepdrix):

0.1139=0.0211x ok good. last step? :d any ideas?

OpenStudy (anonymous):

divide 0.1139 by 0.0211= 5.39 = 5.4

OpenStudy (anonymous):

so 5?

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

Yay! thank you so much! :)

zepdrix (zepdrix):

np :3